OpenStudy (anonymous):

, Find an equation of the tangent line to the graph of x^2+3y^2=4 at the point (1,1)... Please and thank you :)

OpenStudy (anonymous):

I did it!

OpenStudy (anonymous):

Good job!... Wanna let the incredibly confused in on it?

OpenStudy (anonymous):

x^2+3y^2=4 at the point (1,1)... 2x+6yy '=0 y '=-2/6=-1/3 at point (1,1) y-1=-(1/3)(x-1) ans or y=-(1/3)x+ 4/3 ans